the principle that will amount to R 10 000 in 8 months time if the simple interest rate is 9,75% per year is??
Answer by stanbon(75887)

the principle that will amount to R 10 000 in 8 months time if the simple interest rate is 9,75% per year is??
------
A(t) = P(1+r*t)
-------
P = A(t)/(1+r*t)
------
Ans: P = 10,000/(1+0.0975*(2/3)) = 10,000/1.065 = $9389.67
-------------
Cheers,
Stan H.
-------------